Any help with this problem would be great...
What rate of interest, to the nearest tenth of a percent, compounded annually is needed for an investment of $1,500 to grow to $3,300 in four years?
Thanks to all the tutors out there.

Answer by checkley77(12844)

1500(1+r)^4=3300
(1+r)^4=3300/1500
(1+r)^4=2.2
1+r=4thrt(2.2)
1+r=1.21788
r=1.21788-1
r=.21788 ans. for the rate.
Proof:
1500(1+.21788)^4=3300
1500(1.21788)^4=3300
1500*2.2=3300
3300=3300